# Negate or Embrace: On How Misalignment Shapes Multimodal Representation Learning

### v2

@@ 1 Introduction
Modern multimodal learning has achieved remarkable success by jointly modeling information from heterogeneous sources such as vision, language, and audio. In particular, multimodal contrastive learning (MMCL) on paired data has emerged as a dominant strategy for aligning modalities (Radford et al., 2021; Jia et al., 2021; Wu et al., 2022)—notably exemplified by vision-language models like CLIP, which attempt to learns a common representation space by maximizing the similarity of real image-text pairs while minimizing that of incorrect pairs (Radford et al., 2021).
However, one fundamental assumption in multimodal contrastive learning is that the training pairs are completely aligned across modalities, meaning that they convey exactly the same semantic information (Xia et al., 2023; Liu et al., 2024d). This assumption, though convenient, is often violated in real-world scenarios, where multimodal data is inherently noisy or imprecisely paired (Miech et al., 2019; Nakada et al., 2023), e.g., when the text only partially describes the paired image content, or even contains descriptive elements that are irrelevant or misleading, a situation generally referred to as misalignment. For example, in a large-scale video-text dataset, over 50% of the purportedly aligned clip-caption pairs were found to be misaligned (Miech et al., 2019). It is worth noting that the impact of misalignment is primarily concentrated in the text modality, as text captions—often generated based on images—are more susceptible to semantic incompleteness and interpretive variability compared to the typically fixed visual semantic aspects.
@@
In light of this, we offer a theoretical perspective that not only facilitates the understanding of misalignment but also provides insights into real-world applications. Specifically, we formulate the problem via a latent variable model (LVM), which captures the underlying generative process of image-text data with misalignment, as shown in Figure 1. In it, the latent space consists of shared semantic variables representing factors common to both modalities (e.g., object shapes and colors), along with modality-specific subspaces that capture unique variations in images and text.
To model misalignment, we introduce two mechanisms: selection bias and perturbation bias. Both act on the shared semantic information but differ in their effects. Selection bias determines which semantic information is preserved in the text. For example, when describing an object, the text might preserve information about its color ("black") but omit details about its texture or shape. On the other hand, perturbation bias introduces errors, such as changing "black" (correct color) to "red" (incorrect color). To make the proposed LVM adaptable to a wide range of real-world scenarios, we allow for an arbitrary causal structure among the latent semantic variables, providing flexibility in multimodal contexts. Finally, given their differences, we model the modalities with separate generative processes.
@@
We summarize our key contributions as follows, and a detailed discussion of related work is provided in § 6.
  1. 1.
    We propose a latent variable model for multimodal data generation that explicitly captures misalignment through two mechanisms: selection bias and perturbation bias (§ 3).
  2. 2.
    We establish a general identifiability result, showing that MMCL recovers the subset of semantic variables unaffected by these biases, independent of the underlying latent causal structure (§ 4.1).
  3. 3.
    We extend this result to two practical scenarios—tasks requiring common representations and those targeting invariant representations—offering actionable insights into how misalignment should inform real-world applications (§ 4.2).
  4. 4.
    We empirically validate our theoretical findings through extensive experiments on both real-world and synthetic image-text datasets, across diverse misalignment settings, including those with structured latent dependencies (§ 5).
@@ 3.1 A Latent Variable Model Characterizing Misalignment
Let $\mathcal{P}_{+}(\mathbb{I}_{\mathbf{s}})$ denote the set of all non-empty subsets [^2] Without loss of generality, we fix a graded lexicographic order over $\mathcal{P}_{+}(\mathbb{I}_{\mathbf{s}})$. See Clar. B.2 for a clarification. of the index set $\mathbb{I}_{\mathbf{s}}$, defined as $\mathcal{P}_{+}(\mathbb{I}_{\mathbf{s}}):=\mathcal{P}(\mathbb{I}_{\mathbf{s}})\setminus\{\emptyset\}$, where $\mathcal{P}(\cdot)$ denotes the power set. The selection bias $\theta$ is defined as an integer index in the range $\theta\in\{1,\dots,2^{n_{s}}-1\}$, corresponding to a specific non-empty semantic subset $\mathbb{I}_{\theta}\in\mathcal{P}_{+}(\mathbb{I}_{\mathbf{s}})$. The complement $\mathbb{I}_{\theta}^{c}=\mathbb{I}_{\mathbf{s}}\setminus\mathbb{I}_{\theta}$ denotes the omitted semantic subset.
@@
Building upon the previously defined selection bias $\theta$ and perturbation bias $\rho$, we now formalize the text generation process, explicitly capturing the misalignment induced by these two biases. Consider an image $\mathbf{x}$ generated by Eq. 3, with associated semantic variables $\mathbf{s}=(\mathbf{s}_{\mathbb{I}_{\theta}},\mathbf{s}_{\mathbb{I}_{\theta}^{c}})$, where the index set $\mathbb{I}_{\theta}$ is determined by $\theta$. For the corresponding text $\mathbf{t}^{\scriptscriptstyle(\theta)}$, we define the latent variables as $\mathbf{z}_{t^{\scriptscriptstyle(\theta)}}=(\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}},\mathbf{m}_{t})$, where $\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}$ represents the perturbed semantic variables under perturbation bias $\rho$, and $\mathbf{m}_{t}$ denotes the text-specific latent variables. The text generation process is then formalized as:
@@ 3.2 Model Assumptions for Theoretical Analysis
Assumption 3.1 (Latent Variables with Smooth Positive Densities).
@@
Given a selection bias $\theta$ and a perturbation bias $\rho$, consider an image-text pair $(\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)})$ generated by Eq. 3 and Eq. 4, respectively. The conditional distribution $p_{\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}\mid\mathbf{s},\theta,\rho}$ is defined via a randomly sampled perturbation subset $A\subseteq\mathbb{I}_{\rho}$, such that:
@@
Note that Eq. 5 essentially implies that, in each $(\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)})$, only a subset of semantic variables $A$ undergoes perturbations, regardless of the underlying causal structure among latent semantic variables. The rationale is that latent semantic variables can only be modified indirectly by altering observations, rather than through direct intervention, unless the latent causal structure is fully identified. In the text modality, it occurs through the misassignment of certain content words to specific image semantics during the captioning process. Unlike a direct intervention, this misalignment does not propagate to descendant semantic variables, as illustrated in Figure 2.
@@ 4.1 Identifiability of Latent Semantic Variables under Misalignment
Let $(\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)})$ be image-text pairs drawn from the data-generating process described in § 3, where $\mathbf{x}$ is generated according to Eq. 3 and $\mathbf{t}_{\theta}$ is generated by Eq. 4. Suppose that Asms. 3.1 and 3.2 hold. Denote by $\mathbf{s}_{\mathbb{I}^{c}_{\rho}}$ the subset of semantic variables that preserved without bias in the text, and define its dimension as $n=|\mathbb{I}_{\theta}|-|\mathbb{I}_{\rho}|$. Let $\mathbf{f}_{x}:\mathcal{X}\to(0,1)^{n}$ and $\mathbf{f}_{t}:\mathcal{T}^{\scriptscriptstyle(\theta)}\to(0,1)^{n}$ be sufficiently flexible, smooth functions. Then, Minimizing the loss $\mathcal{L}_{\text{SymAlignMaxEnt}}$ in Eq. 2 over samples $(\mathbf{x},\mathbf{t}_{\theta})$ guarantees that $\mathbf{f}_{x}$ and $\mathbf{f}_{t}$ block-identify the semantic variables $\mathbf{s}_{\mathbb{I}^{c}_{\rho}}$ in the sense of Defn. 4.1.
@@ 4.2 Insights into Misalignment for Practice
Let the selection bias be $\theta=2^{n_{s}}-1$ and the perturbation bias be $\rho=1$, such that the full set of semantic variables $\mathbb{I}_{\mathbf{s}}$ is selected, and the perturbable semantic subset is trivial, i.e., $\mathbb{I}_{\rho}=\emptyset$. Then, all semantic variables $\mathbf{s}$ are block-identified via smooth functions $\mathbf{f}_{x}:\mathcal{X}\to(0,1)^{n_{s}}$ and $\mathbf{f}_{t}:\mathcal{T}^{\scriptscriptstyle(\theta)}\to(0,1)^{n_{s}}$, when minimizing $\mathcal{L}_{\text{SymAlignMaxEnt}}$.
@@
Consider an OOD setting in which a subset of semantic variables, $\mathbb{I}_{inv}\subset\mathbb{I}_{\mathbf{s}}$, remains invariant between training and testing environments, while the remaining semantic variables, $\mathbb{I}_{var}=\mathbb{I}_{\mathbf{s}}\setminus\mathbb{I}_{inv}$, undergo distribution shifts. If the union of omitted and perturbable semantic variables under selection bias $\theta$ and perturbation bias $\rho$ coincides with the environment-sensitive subset, i.e., $\mathbb{I}_{var}=\mathbb{I}_{\theta}^{c}\cup\mathbb{I}_{\rho},$ then the invariant semantic variables $\mathbf{s}_{\mathbb{I}_{inv}}$ are block-identified via smooth functions $\mathbf{f}_{x}:\mathcal{X}\to(0,1)^{|\mathbb{I}_{inv}|}$ and $\mathbf{f}_{t}:\mathcal{T}^{\scriptscriptstyle(\theta)}\to(0,1)^{|\mathbb{I}_{inv}|},$ by minimizing $\mathcal{L}_{\text{SymAlignMaxEnt}}$.
@@ 5.1 Numerical Simulation
We synthesize numerical data following the generative process described in § 3. Specifically, we sample modality-specific variables $\mathbf{m}_{x}\sim\mathcal{N}(0,\Sigma_{\mathbf{m}_{x}})$ and $\mathbf{m}_{t}\sim\mathcal{N}(0,\Sigma_{\mathbf{m}_{t}})$, each of dimension 5, along with full semantic variables $\mathbf{s}\sim\mathcal{N}(0,\Sigma_{\mathbf{s}})$, of dimension 10. Potential causal dependencies among these variables are encoded in their covariance matrices $\Sigma_{(\cdot)}$. To simulate various misalignment scenarios, we progressively increase the strength of selection and perturbation biases. For selection bias, we incrementally define selected subsets $\mathbb{I}_{\theta}$ as $\{1\},\dots,[10]$. [^4] For any positive integer $n>1$, we use the notation $[n]$ to denote the set $\{1,2,\dots,n\}$. For perturbation bias, we similarly use increasing subsets of $\mathbb{I}_{\rho}$, ranging from the empty set to $[9]$, applying additive Gaussian noise with a probability 0.75 for each semantic dimension $i\in\mathbb{I}_{\rho}$. These biases are applied separately: when analyzing selection bias, we set $\mathbb{I}_{\rho}=\emptyset$; conversely, when analyzing perturbation bias, we fix $\mathbb{I}_{\theta}=[9]$. Generation functions $\mathbf{g}_{x}$ and $\mathbf{g}_{t_{\theta}}$ are instantiated as randomly initialized invertible MLPs. Two modality-specific MLP encoders are trained for 100,000 steps using the $\mathcal{L}_{\text{SymAlignMaxEnt}}$ loss defined in Eq. 2, setting the representation dimension equal to the unbiased semantic dimensions. Further details on the experimental setup are provided in § C.1.
We conduct two main experiments. (i) Identification of semantics: We predict each dimension of the true semantic variables from the learned representations using a lightweight MLP, reporting the predictive $R^{2}$ score on holdout data. Results are averaged over three random seeds for each setting. (ii) Downstream performance: We evaluate the pretrained representations obtained under various bias conditions on downstream tasks. Specifically, we construct four regression task labels $y_{1},y_{2},y_{3}$, and $y_{4}$ by applying complex nonlinear functions to subsets of the semantic variables: $[3]$, $[5]$, $[7]$, and $[9]$, respectively. For classification, we binarize $y_{2}$ to produce binary labels. To evaluate the OOD generalization performance of the trained classifiers, we introduce a distribution shift in the observations $\mathbf{x}$ by applying a heavy-tailed transformation to dimensions $\{9,10\}$. Further details on the task design are provided in § C.1.
@@ 5.2 MPI3d-Complex Dataset
Training is performed using the multimodal contrastive loss $\mathcal{L}_{\text{MMCL}}$ defined in Eq. 1, following the formulation in (Daunhawer et al., 2022). Performance is evaluated using the average Matthews Correlation Coefficient (MCC) across three random seeds, based on the prediction of all image latent variables from the learned image and text representations using linear and nonlinear MLP decoders, respectively. Further implementation details—including dataset statistics, bias configurations, text generation procedures, encoder architectures, and training hyperparameters—are provided in § D.1.

### v3

@@ 1 Introduction
Modern multimodal learning has achieved remarkable success by jointly modeling information from heterogeneous sources such as vision, language, and audio. In particular, multimodal contrastive learning (MMCL) on paired data has emerged as a dominant strategy for aligning modalities (Radford et al., 2021; Jia et al., 2021; Wu et al., 2022)—notably exemplified by vision-language models like CLIP, which attempt to learns a common representation space by maximizing the similarity of real image-text pairs while minimizing that of incorrect pairs (Radford et al., 2021). However, one fundamental assumption in MMCL is that the training pairs are completely aligned across modalities, meaning that they convey exactly the same semantic information (Xia et al., 2023; Liu et al., 2024d). This assumption, though convenient, is often violated in real-world scenarios, where multimodal data is inherently noisy or imprecisely paired (Miech et al., 2019; Nakada et al., 2023), e.g., when the text only partially describes the paired image content, or even contains descriptive elements that are irrelevant or misleading, a situation generally referred to as misalignment. For example, in a large-scale video-text dataset, over 50% of the purportedly aligned clip-caption pairs were found to be misaligned (Miech et al., 2019). It is worth noting that the impact of misalignment is primarily concentrated in the text modality, as text captions—often generated based on images—are more susceptible to semantic incompleteness and interpretive variability compared to the typically fixed visual semantic aspects.
@@
In light of this, we offer a theoretical perspective that not only facilitates the understanding of misalignment but also provides insights into real-world applications. Specifically, we formulate the problem via a latent variable model (LVM), which depicts the underlying generative process of image-text data with misalignment, as shown in Figure 1. In it, the latent space consists of shared semantic variables representing factors common to both modalities (e.g., object shapes and colors), along with modality-specific subspaces that capture unique variations in images and text. To model misalignment, we introduce two mechanisms: selection bias and perturbation bias. Both act on the shared semantic information but differ in their effects. Selection bias determines which semantic information is preserved in the text. For example, when describing an object, the text might preserve information about its color ("black") but omit details about its texture or shape. On the other hand, perturbation bias introduces errors, such as changing "black" (correct color) to "red" (incorrect color). To make the proposed LVM adaptable to a wide range of real-world scenarios, we allow for an arbitrary causal structure among the latent semantic variables, providing flexibility in multimodal contexts. Finally, given their differences, we model the modalities with separate generative processes.
@@
We summarize our key contributions as follows, and a detailed discussion of related work is provided in § 6. (i) We propose a latent variable model for multimodal data generation that explicitly captures misalignment through two mechanisms: selection bias and perturbation bias (§ 3). (ii) We establish a general identifiability result, showing that MMCL recovers the subset of semantic variables unaffected by these biases, independent of the underlying latent causal structure (§ 4.1). (iii) We extend this result to two practical scenarios—tasks requiring common representations and those targeting invariant representations—offering actionable insights into how misalignment should inform real-world applications (§ 4.2). (iv) We empirically validate our theoretical findings through extensive experiments on both real-world and synthetic image-text datasets, across diverse misalignment settings, including those with structured latent dependencies (§ 5).
@@ 3.1 A Latent Variable Model Characterizing Misalignment
Let $\mathcal{P}_{+}(\mathbb{I}_{\mathbf{s}})$ denote the set of all non-empty subsets [^2] Without loss of generality, we fix a graded lexicographic order over $\mathcal{P}_{+}(\mathbb{I}_{\mathbf{s}})$. See Clar. B.2 for a clarification. of the index set $\mathbb{I}_{\mathbf{s}}$, defined as $\mathcal{P}_{+}(\mathbb{I}_{\mathbf{s}}):=\mathcal{P}(\mathbb{I}_{\mathbf{s}})\setminus\{\emptyset\}$, where $\mathcal{P}(\cdot)$ denotes the power set. The selection bias $\theta$ is defined as an integer index in the range $\{1,\dots,2^{n_{s}}-1\}$, corresponding to a specific non-empty semantic subset $\mathbb{I}_{\theta}\in\mathcal{P}_{+}(\mathbb{I}_{\mathbf{s}})$. The complement $\mathbb{I}_{\theta}^{c}=\mathbb{I}_{\mathbf{s}}\setminus\mathbb{I}_{\theta}$ denotes the omitted semantic subset.
@@
Building upon the previously defined selection bias $\theta$ and perturbation bias $\rho$, we now formalize the text generation process, explicitly characterizing the misalignment induced by these two biases. Consider an image $\mathbf{x}$ generated by Eq. 3, with associated semantic variables $\mathbf{s}=(\mathbf{s}_{\mathbb{I}_{\theta}},\mathbf{s}_{\mathbb{I}_{\theta}^{c}})$, where the index set $\mathbb{I}_{\theta}$ is determined by $\theta$. For the corresponding text $\mathbf{t}^{\scriptscriptstyle(\theta)}$, we define the latent variables as $\mathbf{z}_{t^{\scriptscriptstyle(\theta)}}=(\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}},\mathbf{m}_{t})$, where $\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}$ represents the perturbed semantic variables under perturbation bias $\rho$, and $\mathbf{m}_{t}$ denotes the text-specific latent variables. The text generation process is then formalized as:
@@ 3.2 Model Assumptions for Theoretical Analysis
Assumption 3.1 (Latent Variables with Continuous Positive Densities).
@@
Given a selection bias $\theta$ and a perturbation bias $\rho$, consider a specific image-text pair $(\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)})$ generated by Eq. 3 and Eq. 4, respectively. The conditional distribution $p_{\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}\mid\mathbf{s},\theta,\rho}$ is defined via a randomly sampled perturbation subset $A\subseteq\mathbb{I}_{\rho}$, such that:
@@
Note that Eq. 5 essentially implies that, in each $(\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)})$, only a subset $A$ of $\mathbb{I}_{\rho}$ undergoes perturbations, regardless of the underlying causal structure among latent semantic variables. The rationale is that latent semantic variables can only be modified indirectly by altering observations, rather than through direct intervention, unless the latent causal structure is fully identified. In the text modality, it occurs through the misassignment of certain content words to specific image semantics during the captioning process. Unlike a direct intervention, this misalignment does not propagate to descendant semantic variables, as illustrated in Figure 2.
@@ 4.1 Identifiability of Latent Semantic Variables under Misalignment
Let $(\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)})$ be an image-text pair drawn from the data-generating process described in § 3, where $\mathbf{x}$ is generated according to Eq. 3 and $\mathbf{t}_{\theta}$ is generated by Eq. 4. Suppose that Asms. 3.1 and 3.2 hold. Denote by $\mathbf{s}_{\mathbb{I}^{c}_{\rho}}$ the subset of semantic variables that preserved without bias in the text, and define its dimension as $n=|\mathbb{I}_{\theta}|-|\mathbb{I}_{\rho}|$. Let $\mathbf{f}_{x}:\mathcal{X}\to(0,1)^{n}$ and $\mathbf{f}_{t}:\mathcal{T}^{\scriptscriptstyle(\theta)}\to(0,1)^{n}$ be sufficiently flexible, smooth functions. Then, minimizing the loss $\mathcal{L}_{\text{SymAlignMaxEnt}}$ in Eq. 2 over samples $(\mathbf{x},\mathbf{t}_{\theta})$ guarantees that $\mathbf{f}_{x}$ and $\mathbf{f}_{t}$ block-identify the semantic variables $\mathbf{s}_{\mathbb{I}^{c}_{\rho}}$ in the sense of Defn. 4.1.
@@ 4.2 Insights into Misalignment for Practice
Assume that Asms. 3.1 and 3.2 hold. Let the selection bias be $\theta=2^{n_{s}}-1$ and the perturbation bias be $\rho=1$, such that the full set of semantic variables $\mathbb{I}_{\mathbf{s}}$ is selected, and the perturbable semantic subset is trivial, i.e., $\mathbb{I}_{\rho}=\emptyset$. Then, all semantic variables $\mathbf{s}$ are block-identified via smooth functions $\mathbf{f}_{x}:\mathcal{X}\to(0,1)^{n_{s}}$ and $\mathbf{f}_{t}:\mathcal{T}^{\scriptscriptstyle(\theta)}\to(0,1)^{n_{s}}$, when minimizing $\mathcal{L}_{\text{SymAlignMaxEnt}}$.
@@
Assume that Asms. 3.1 and 3.2 hold. Consider an OOD setting in which a subset of semantic variables, $\mathbb{I}_{inv}\subset\mathbb{I}_{\mathbf{s}}$, remains invariant between training and testing environments, while the remaining semantic variables, $\mathbb{I}_{var}=\mathbb{I}_{\mathbf{s}}\setminus\mathbb{I}_{inv}$, undergo distribution shifts. If the union of omitted and perturbable semantic variables under selection bias $\theta$ and perturbation bias $\rho$ coincides with the environment-sensitive subset, i.e., $\mathbb{I}_{var}=\mathbb{I}_{\theta}^{c}\cup\mathbb{I}_{\rho},$ then the invariant semantic variables $\mathbf{s}_{\mathbb{I}_{inv}}$ are block-identified via smooth functions $\mathbf{f}_{x}:\mathcal{X}\to(0,1)^{|\mathbb{I}_{inv}|}$ and $\mathbf{f}_{t}:\mathcal{T}^{\scriptscriptstyle(\theta)}\to(0,1)^{|\mathbb{I}_{inv}|},$ by minimizing $\mathcal{L}_{\text{SymAlignMaxEnt}}$.
@@ 5.1 Numerical Simulation
We simulate data following the generative process in § 3. Semantic variables $\mathbf{s}\sim\mathcal{N}(0,\Sigma_{\mathbf{s}})$ (dim. 10) and modality-specific variables $\mathbf{m}_{x}\sim\mathcal{N}(0,\Sigma_{\mathbf{m}_{x}}),\mathbf{m}_{t}\sim\mathcal{N}(0,\Sigma_{\mathbf{m}_{t}})$ (dim. 5) encode potential causal dependencies via their respective covariances $\Sigma_{(\cdot)}$. To study misalignment, we independently vary: (i) Selection bias: controlling $\mathbb{I}_{\theta}=\{1\},\dots,[10]$ [^4] For any positive integer $n>1$, we use the notation $[n]$ to denote the set $\{1,2,\dots,n\}$., and (ii) Perturbation bias: defining $\mathbb{I}_{\rho}=\emptyset,\dots,[9]$, with Gaussian noise added to each $i\in\mathbb{I}_{\rho}$ with probability 0.75. In isolation, we fix $\mathbb{I}_{\rho}=\emptyset$ when varying $\mathbb{I}_{\theta}$, and $\mathbb{I}_{\theta}=[10]$ when varing $\mathbb{I}_{\rho}$. Image and text data are generated using randomly initialized invertible MLPs $\mathbf{g}_{x},\mathbf{g}_{t^{\scriptscriptstyle\theta}}$. We train two modality-specific MLP encoders for 100,000 steps using the $\mathcal{L}_{\text{SymAlignMaxEnt}}$ loss (Eq. 2), with embedding dimension matching the number of unbiased semantics. See § C.1 for further details.
We conduct two experiments. (i) Semantic identification: A lightweight MLP regressor is trained to recover each ground-truth semantic dimension from learned embeddings; we report mean $R^{2}$ over three seeds. (ii) Downstream performance: We evaluate embeddings on four tasks with targets $y_{1},y_{2},y_{3},y_{4}$, defined as nonlinear functions over semantic subsets [3], [5], [7], [9]. We additionally binarize $y_{2}$ for classification, and apply a heavy-tailed shift to dimensions $\{9,10\}$ of semantic variables to evaluate OOD generalization. See § C.1 for full task design.
@@ 5.2 MPI3D-Complex Dataset
The MPI3D-Complex dataset (Gondal et al., 2019) contains real-world images annotated with seven mutually independent, discrete ground-truth factors: object color (color), shape (shape), size (size), camera height (cam.), background color (back.), horizontal position (hori.), and vertical position (vert.). We treat hori. and vert. as image-specific factors, and the remaining five as semantic variables. See Figure 5 for example images from the MPI3D-Complex dataset.
@@
Training is performed using the loss $\mathcal{L}_{\text{MMCL}}$ defined in Eq. 1, following the formulation in (Daunhawer et al., 2022). Performance is evaluated using the average Matthews Correlation Coefficient (MCC) across three random seeds, based on the prediction of all image latent variables from the learned image and text representations using linear and nonlinear MLP decoders, respectively. Full details on dataset attributes, text generation, bias configurations, and architectures are provided in § D.1.